# reference.bib

### v1

@@ 7. Differential Operators
In this section, we prepare the ground for the theta-correspondence. In the same fashion to Krieg in [krieg1991], Gritsenko in [gritsenko] and Raghavan and Sengupta in [raghavan], we will need to apply some differential operators to the theta series first, so that the integral converges. Our first step is to make $\Theta$ invariant under the action of $\hbox{Sp}_{2}(\mathbb{Z})$, up to the character $\chi_{S}$. This is essential because the differential operators that will eliminate the terms of the theta series that cause divergence are $\hbox{Sp}_{2}(\mathbb{R})$-invariant. From now on, we assume that $4\mid n$.
@@
$\displaystyle\delta_{k}:=\det(Y)^{-k+\frac{m-1}{2}}\det(\partial_{Z})\det(Y)^{k-\frac{m-1}{2}},$
@@
$\delta_{k}^{(r)}:=\delta_{k+2r-2}\cdots\delta_{k+2}\delta_{k}.$
@@
For any dimension $m\geq 1$, we consider the algebras $\mathbb{D}(\mathbb{H}_{m})$ and $\mathbb{D}(\mathcal{P}_{m})$ of invariant differential operators with respect to $\hbox{Sp}_{m}(\mathbb{R})$ and $\hbox{GL}_{m}(\mathbb{R})$. Here, $\mathcal{P}_{m}$ denotes the symmetric space $\hbox{GL}_{m}(\mathbb{R})/\hbox{O}_{m}(\mathbb{R})$ and can be identified with the set
@@
by means of the action of $\hbox{GL}_{m}(\mathbb{R})$ on $\mathcal{P}_{m}$ given by
@@
which sends $D\longmapsto\phi^{-1}\circ D\circ\phi$. Then, by [deitmar_krieg, Theorem 1.1], we obtain that $\phi^{*}$ is an injective algebra homomorphism.
@@
By [deitmar_krieg, Theorem 1.2], we have that $D_{m+2-l,m}D_{l,m}$ belongs in the image of $\phi^{*}$. Define then
@@
We now have the following Lemma regarding the behaviour of $\Theta$ under the action of the Maass-Shimura operator.
@@
Let $k=-n/2$ and $r=n/4$, as before. For any $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$, we have
$\delta_{k}^{(r)}\left[(\det Y)^{\frac{n+2}{2}}\theta_{\ell}(Z,W)\right]=f_{\ell}(S_{1}[\ell],R_{W}[\ell],Y)\theta_{\ell}(Z,W),$
for some polynomial $f_{\ell}$ in the entries of $S_{1}[\ell]\textup{, }R_{W}[\ell]\textup{, }Y$. However, if $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ is such that $S_{1}[\ell]=0$, we have
$\delta_{k}^{(r)}\left[(\det Y)^{\frac{n+2}{2}}\theta_{\ell}(Z,W)\right]=(\det Y)^{-r}p\left(\det(R_{W}[\ell]Y),\textup{tr}\left(R_{W}[\ell]Y\right)\right)e^{-\pi\textup{tr}\left(R_{W}\left[\ell\right]Y\right)},$
where $p=p(U,V)\in\mathbb{R}[U,V]$ is a polynomial in two variables, and its coefficients do not depend on $\ell$.
@@
We first have by [satoh1986differential, Lemma 1.1]
@@
But now by [satoh1986differential, Proposition 1.2, (a)], $\delta_{1}^{(r)}\theta_{\ell}(Z,W)$ is a $\mathbb{Z}[1/2]$-linear combination of functions of the form
$\det(Y)^{-b}\sigma^{c}\left(\det(\partial_{Z})\right)^{d}\theta_{\ell}(Z,W),$
with $0\leq b\leq c\leq r,0\leq d\leq r$ and $b+d=r$. But as in Proposition 7.1, we have
$\frac{\partial}{\partial z_{1}}\theta_{\ell}(Z,W)=A_{\ell}\theta_{\ell}(Z,W),\textup{ }\frac{\partial}{\partial z_{2}}\theta_{\ell}(Z,W)=B_{\ell}\theta_{\ell}(Z,W),\textup{ }\frac{\partial}{\partial z_{3}}\theta_{\ell}(Z,W)=2C_{\ell}\theta_{\ell}(Z,W),$
with $A_{\ell},B_{\ell},C_{\ell}$ as defined there. The first part of the lemma then follows by the form of the operators $\det(\partial_{Z})$ and $\sigma$. In the case $S_{1}[\ell]=0$, we can write each term above as
$\left(-\frac{1}{4}\pi^{2}\right)^{d}\det(Y)^{-r}[(\det Y)^{r-b}\det(R_{W}[\ell])^{d}\sigma^{c}\theta_{\ell}(Z,W)]=\left(-\frac{1}{4}\pi^{2}\right)^{d}\det(Y)^{-r}[\det(R_{W}[\ell]Y)^{d}\sigma^{c}\theta_{\ell}(Z,W)],$
because $\det(\partial_{Z})\theta_{\ell}=-\pi^{2}\det(R_{W}[\ell])\theta_{\ell}/4$ and $r-b=d$. The result now follows by observing that $\sigma^{c}\theta_{\ell}$ is a polynomial with real coefficients in $\textup{tr}\left(R_{W}[\ell]Y\right)$ (by an inductive argument, using the relations derived in Proposition 7.1). ∎
@@
What is more, given $l\in\mathbb{R}\textup{, }\epsilon>0$, and a compact subset $\mathcal{C}$ of $\mathcal{H}_{S}$, there is a constant $C>0$ such that
holds for all $W\in\mathcal{C}$ and $Z=X+iY\in\mathbb{H}_{2}$, with $Y\geq\epsilon I$.
@@
Fix $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ such that rank <$\textup{rank }\ell<2$. We may assume that
@@
for some $x,y\in\mathbb{R}$. This is true because we can find $U\in\textup{SL}_{2}(\mathbb{R})$ such that $\ell U=\begin{pmatrix}a&0\end{pmatrix}$ for some $a\in\mathbb{R}^{n+4}$. But then $U_{1}:=\begin{pmatrix}U&0\\ 0&(U^{t})^{-1}\end{pmatrix}\in\textup{Sp}_{2}(\mathbb{R})$ and the action of this matrix on $\mathbb{H}_{2}$ is $Z\longmapsto UZU^{t}$. Since $R$ is $\textup{Sp}_{2}$-invariant, we can change variables $Z\longmapsto UZU^{t}$. But
@@
so we can replace $\ell$ with $\ell U$, which will then give the form of $S_{1}[\ell]$ wanted. Similarly for $R_{W}[\ell]$.
We now have
@@
by [satoh1986differential, Lemma 1.1] (we remind here that $k=-n/2$). Define also the operator $\sigma\in\mathbb{D}(\mathbb{H}_{2})$ by
$\sigma:=i\sum_{j=1}^{3}y_{j}\frac{\partial}{\partial z_{j}}.$
By [satoh1986differential, Proposition 1.2, (a)], we have that the quantity $\delta_{1}^{(r)}\theta_{\ell}(Z,W)$ will be a $\mathbb{Z}[1/2]$-linear combination of functions of the form
@@
with $0\leq b\leq c\leq r,0\leq d\leq r$ and $b+d=r$. Now, we compute
$\frac{\partial}{\partial z_{1}}\theta_{\ell}(Z,W)=A_{\ell}\theta_{\ell}(Z,W),$
where $A_{\ell}=\frac{1}{2}\pi i\left(S_{1}[l]+R_{W}[l]\right)$. Here, we write $\ell=\begin{pmatrix}l&m\end{pmatrix}$, with $l,m\in\mathbb{Z}^{n+4}$. Similarly,
$\frac{\partial}{\partial z_{2}}\theta_{\ell}(Z,W)=B_{\ell}\theta_{\ell}(Z,W),\textup{ }\frac{\partial}{\partial z_{3}}\theta_{\ell}(Z,W)=2C_{\ell}\theta_{\ell}(Z,W),$
where $B_{\ell}=\frac{1}{2}\pi i\left(S_{1}[m]+R_{W}[m]\right)$ and $C_{\ell}=\frac{1}{2}\pi i\left(l^{t}S_{1}m+l^{t}R_{W}m\right)$. But in the case rank <$\textup{rank }\ell<2$, we have that $B_{\ell}=C_{\ell}=0$ because of the form of $S_{1}[\ell]$ and $R_{W}[\ell]$. Hence,
@@
So, we only need to consider the case $b=c=r,d=0$. But
$\sigma\theta_{\ell}(Z,W)=iA_{\ell}y_{1}\theta_{\ell}(Z,W),$
and therefore, by an inductive argument, $\sigma^{r}\theta_{\ell}$ depends only on $y_{1}$ and $\theta_{\ell}$. Consider now the map
@@
$\theta_{\ell}(Z,W)=e\left(\frac{1}{2}\left(xX_{1}^{\prime}+iyY_{1}^{\prime}\right)\right),$
@@
is independent of $a,b$ and so belongs in the image of $\tau$. But by [deitmar_krieg, Proposition 1], $\tau^{-1}\circ R\circ\tau$ is a simple tensor in $\mathbb{D}(\mathbb{H}_{1}\times\mathbb{H}_{1})$, so the problem is reduced in the one-dimensional case and $S_{1}[\ell]=R_{W}[\ell]=0$, i.e. suffices to show
@@
where now we have $\delta_{k}=\frac{k}{2iy}+\frac{\partial}{\partial z}$ in the one-dimensional case. But we have
$\delta_{k}^{(r)}y^{\frac{n+2}{2}}=\textup{(const)}\times y^{n/4+1}$
and since by [bocherer, page 807] we have that in the one-dimensional case
$R(1,2+n/4)=4y^{2}\frac{\partial}{\partial{Z}}\frac{\partial}{\partial{\overline{Z}}}-\frac{n}{4}\left(\frac{n}{4}+1\right),$
the result follows.
The final part of the Proposition follows in the exact same way as in [deitmar_krieg, Proposition 2.1(b)], using also the first part of Lemma 7.1. ∎
@@ 8. Theta Correspondence
In this section, we finally give the theta correspondence between the Klingen Eisenstein series of $\hbox{SO}(2,n)$ and the Siegel Eisenstein series for $\hbox{Sp}_{2}$. We start with the following definition:
@@
Let $\chi$ denote the character of Theorem 6.1. Let also
$P_{2,0}:=\left\{\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\hbox{Sp}_{2}(\mathbb{Z})\mid C=0\right\}$
denote the Siegel parabolic subgroup of $\hbox{Sp}_{2}$. Notice that $P_{2,0}\cap\Gamma_{0}^{(2)}(q)=P_{2,0}$. For $s\in\mathbb{C}$ with $\hbox{Re}(s)>3/2$, we then define the Siegel Eisenstein series with respect to $P_{2,0}$ and with character $\chi$ as
$\widetilde{E}(Z,\chi,s):=\sum_{\gamma\in P_{2,0}\backslash\Gamma_{0}^{(2)}(q)}\overline{\chi(\gamma)}(\det(\textup{Im}(\gamma Z)))^{s}.$
If $\delta\in P_{2,0}$, we have $\det\delta=\pm 1$, so $\chi(\delta)=\psi(\det\delta)=1$ by Proposition 6.1 and so the Eisenstein series is well-defined. This Eisenstein series has a meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=3/2$, with constant non-zero residue (cf. [feit, Theorem 9.2]). Finally, using the fact that $\chi$ is a character, we have
@@
for all $\gamma\in\Gamma_{0}^{(2)}(q)$. The main Theorem of the paper can be now stated as follows:
@@
Let $S$ have rank $n$, with $4\mid n$ and be such that #$\#\mathcal{C}^{1}(\Gamma_{S})=1$. Let also $k=-n/2$ and $r=n/4$, as before. Define
$\phi_{2}(t):=t\left(t-\frac{1}{2}\right),\textup{ }\xi(s):=\pi^{-s/2}\Gamma(s/2)\zeta(s),$
for $t\in\mathbb{C}$. We then have for $\hbox{Re}(s)>n+1$
@@
where $\gamma_{S}(s)=(-4)^{r}\phi_{2}(s/2-2r)\phi_{2}(s/2)\prod_{j=1}^{r}\phi_{2}(s/2-2j+3/2)$.
In particular, $E(Z,s)$ has a meromorphic continuation to $\mathbb{C}$, due to the one of $\widetilde{E}(Z,\chi,s)$.
@@
Let $I:=[\textup{Sp}_{2}(\mathbb{Z}):\Gamma_{0}^{(2)}(q)]\left\langle\widetilde{E}(Z,\chi,(s+1)/2-r),R[\delta_{k}^{(r)}\Theta](Z,W)\right\rangle_{\Gamma_{0}^{(2)}(q)}$. First of all, this integral is well-defined because of the second part of Proposition 7.1. We then have
[…]

### v2

@@ 7. Differential Operators
In this Section, we prepare the ground for the theta-correspondence. In the same fashion as Krieg in [krieg1991], Gritsenko in [gritsenko], and Raghavan and Sengupta in [raghavan], we need to apply some differential operators to the theta series first, so that the integral converges. Our first step is to make $\Theta$ invariant under the action of $\textup{Sp}_{2}(\mathbb{Z})$, up to the character $\chi_{S}$. This is essential because the differential operators that will eliminate the terms of the theta series that cause divergence are $\textup{Sp}_{2}(\mathbb{R})$-invariant. From now on, we assume that $4\mid n$.
@@
(11) $\delta_{k}:=\det(Y)^{-k+\frac{m-1}{2}}\det(\partial_{Z})\det(Y)^{k-\frac{m-1}{2}},$
@@
(12) $\delta_{k}^{(r)}:=\delta_{k+2r-2}\cdots\delta_{k+2}\delta_{k}.$
@@
For any dimension $m\geq 1$, we consider the algebras $\mathbb{D}(\mathbb{H}_{m})$ and $\mathbb{D}(\mathcal{P}_{m})$ of invariant differential operators with respect to $\textup{Sp}_{m}(\mathbb{R})$ and $\textup{GL}_{m}(\mathbb{R})$. Here, $\mathcal{P}_{m}$ denotes the symmetric space $\textup{GL}_{m}(\mathbb{R})/\textup{O}_{m}(\mathbb{R})$ and can be identified with the set
@@
by means of the action of $\textup{GL}_{m}(\mathbb{R})$ on $\mathcal{P}_{m}$ given by
@@
which sends $D\longmapsto\phi^{-1}\circ D\circ\phi$. Then, from [deitmar_krieg, Theorem 1.1], we obtain that $\phi^{*}$ is an injective algebra homomorphism.
@@
From [deitmar_krieg, Theorem 1.2], we have that $D_{m+2-l,m}D_{l,m}$ belongs in the image of $\phi^{*}$. Define then
@@
We first have the following Lemma regarding the behaviour of $\Theta$ under the action of the Maass-Shimura operator.
@@
Let $k=-n/2$ and $r=n/4$. For any $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$, we have
(13) $\delta_{k}^{(r)}\left[(\det Y)^{\frac{n+2}{2}}\theta_{\ell}(Z,W)\right]=(\det Y)^{1+r}p(\det((S_{1}+R_{W})[\ell]Y),\textup{tr}((S_{1}+R_{W})[\ell]Y))\theta_{\ell}(Z,W),$
where $p=p(U,V)\in\mathbb{R}[U,V]$ is a polynomial in two variables, and its coefficients do not depend on $\ell$. In particular, if $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ is such that $S_{1}[\ell]=0$, this becomes
(14) $\delta_{k}^{(r)}\left[(\det Y)^{\frac{n+2}{2}}\theta_{\ell}(Z,W)\right]=(\det Y)^{1+r}p\left(\det(R_{W}[\ell]Y),\textup{tr}\left(R_{W}[\ell]Y\right)\right)e^{-\pi\textup{tr}\left(R_{W}\left[\ell\right]Y\right)}.$
@@
For any $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$, we have by [satoh1986differential, Lemma 1.1]
@@
Define now the operator $\sigma\in\mathbb{D}(\mathbb{H}_{2})$ by
(15) $\sigma:=i\sum_{j=1}^{3}y_{j}\frac{\partial}{\partial z_{j}}.$
By [satoh1986differential, Proposition 1.2, (a)], $\delta_{1}^{(r)}\theta_{\ell}(Z,W)$ is a $\mathbb{Z}[1/2]$-linear combination of functions of the form
(16) $\det(Y)^{-b}\sigma^{c}\left(\det(\partial_{Z})\right)^{d}\theta_{\ell}(Z,W),$
where $b,c,d$ are integers with $0\leq b\leq c\leq r,\ 0\leq d\leq r$ and $b+d=r$. Note that the additional requirements stated in [satoh1986differential, Proposition 1.2] are not needed for the proof of this part. We compute
(17) $\frac{\partial}{\partial z_{1}}\theta_{\ell}(Z,W)=A_{\ell}\theta_{\ell}(Z,W),\textup{ }\frac{\partial}{\partial z_{2}}\theta_{\ell}(Z,W)=B_{\ell}\theta_{\ell}(Z,W),\ \frac{\partial}{\partial z_{3}}\theta_{\ell}(Z,W)=2C_{\ell}\theta_{\ell}(Z,W),$
where for $\ell=\begin{pmatrix}l&m\end{pmatrix}$, with $l,m\in\mathbb{Z}^{n+4}$ we have
${A_{\ell}=\frac{1}{2}\pi i\left(S_{1}[l]+R_{W}[l]\right)},\ \displaystyle{B_{\ell}=\frac{1}{2}\pi i\left(S_{1}[m]+R_{W}[m]\right)},\ \displaystyle{C_{\ell}=\frac{1}{2}\pi i\left(l^{t}S_{1}m+l^{t}R_{W}m\right)}.$
We now set $\displaystyle{T_{\ell}:=\frac{1}{2}\pi i(S_{1}+R_{W})[\ell]}$ and write $T_{\ell}=\begin{pmatrix}t_{1}&t_{3}/2\\ t_{3}/2&t_{2}\end{pmatrix}$, with $t_{1}=A_{\ell},t_{2}=B_{\ell}$ and $t_{3}=2C_{\ell}$.
We first compute $\det(\partial_{Z})\theta_{\ell}=\det(T_{\ell})\theta_{\ell}$. Moreover, if $B:=\textup{tr}(T_{\ell}Y)$, we claim that for any $0\leq c\leq r$, $\sigma^{c}\theta_{\ell}=f_{c}(B)\theta_{\ell}$ for some polynomial $f$ of degree $c$. We show this by induction. Note that for $j=1,2,3$, we have $\partial B/\partial z_{j}=-it_{j}/2$.
If $c=0$, then the claim is clear with $f_{0}(B)=1$. If now $\sigma^{c}\theta_{\ell}=f_{c}(B)\theta_{\ell}$, we have
$\sigma^{c+1}\theta_{\ell}=i\sum_{j=1}^{3}y_{j}\left[f_{c}^{\prime}(B)\left(\frac{-it_{j}}{2}\right)\theta_{\ell}+f_{c}(B)t_{j}\theta_{\ell}\right]=B\left(\frac{1}{2}f_{c}^{\prime}(B)+if_{c}(B)\right)\theta_{\ell},$
from which the claim follows with $\displaystyle{f_{c+1}(B)=B\left(\frac{1}{2}f_{c}^{\prime}(B)+if_{c}(B)\right)}$ of degree $c+1$.
Now, each term in (16) can be written as
$(\det Y)^{-b}\sigma^{c}\det(\partial_{Z})^{d}\theta_{\ell}=(\det Y)^{-b}\sigma^{c}\det(T_{\ell})^{d}\theta_{\ell}=(\det Y)^{-r}[\det(T_{\ell}Y)^{d}\sigma^{c}\theta_{\ell}]=\\ =(\det Y)^{-r}[\det(T_{\ell}Y)^{d}f_{c}(\textup{tr}(T_{\ell}Y))\theta_{\ell}],$
because $r-b=d$. Equation (13) now follows after absorbing the $\pi i/2$ factor of $T_{\ell}$ and observing (by induction) that $f_{c}(B)$ has purely imaginary coefficients in the odd powers of $B$ and real coefficients in the even powers. Hence $p$ will have real coefficients. Equation (14) now follows immediately from (13). ∎
@@
Due to Maass in [maass] and Shimura in [shimura_differential], $R$ can be written in the form
(18) $R=u(H_{1},H_{2})$
where $H_{1},H_{2}$ are some “generalised” Laplacians and $u\in\mathbb{C}[X,Y]$ a polynomial (see also [yang, Example 3.2] and [bump_choie, Proposition 6] for an explicit description of $H_{1},H_{2}$). Therefore, because $\Theta$ is absolutely and uniformly convergent on compact subsets of $\mathbb{H}_{2}$, we can apply the differential operators term by term. Hence, it suffices to show that for any $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ with $\textup{rank }\ell<2$, we have
$R\left[\delta_{k}^{(r)}\left[(\det Y)^{\frac{n+2}{2}}e^{\pi i\textup{tr}(S_{1}[\ell]X+iR_{W}[\ell]Y)}\right]\right]=0.$
Fix now $\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})$ such that $\textup{rank }\ell<2$. We may assume that
@@
for some $x,y\in\mathbb{R}$. This is true because we can find $U\in\textup{SL}_{2}(\mathbb{R})$ such that $\ell U=\begin{pmatrix}a&0\end{pmatrix}$ for some $a\in\mathbb{R}^{n+4}$. But then
$\widetilde{U}:=\begin{pmatrix}U&0\\ 0&(U^{t})^{-1}\end{pmatrix}\in\textup{Sp}_{2}(\mathbb{R})$
and the action of this matrix on $\mathbb{H}_{2}$ is $Z\longmapsto UZU^{t}$. Since $R$ is $\textup{Sp}_{2}$-invariant and $j(\widetilde{U},Z)=1$, $\psi_{S}(\det U)=1$ (so the action $Z\longmapsto\widetilde{U}\langle Z\rangle$ does not change $\delta_{k}^{(r)}\Theta$), we can change variables $Z\longmapsto UZU^{t}$. But $\det(UYU^{t})=\det(Y)$ and
@@
so we can replace $\ell$ with $\ell U$, which will then give the form of $S_{1}[\ell]$ wanted. Similarly for $R_{W}[\ell]$. By [satoh1986differential, Lemma 1.1], we have (we remind here that $k=-n/2$)
@@
and by [satoh1986differential, Proposition 1.2, (a)], we have that the quantity $\delta_{1}^{(r)}\theta_{\ell}(Z,W)$ will be a $\mathbb{Z}[1/2]$-linear combination of functions of the form
@@
where $b,c,d$ are integers with $0\leq b\leq c\leq r,\ 0\leq d\leq r$ and $b+d=r$ (the operator $\sigma$ here is as in (15)). Now, in the case $\textup{rank }\ell<2$, we have that $B_{\ell}=C_{\ell}=0$ in (17) because of the form of $S_{1}[\ell]$ and $R_{W}[\ell]$. Hence,
@@
So, we only need to consider the case $b=c=r,\ d=0$. But from the proof of Lemma 7.1, we have $\sigma^{r}\theta_{\ell}(Z,W)=f_{r}(B)\theta_{\ell}(Z,W)$, with $B=\textup{tr}(T_{\ell}Y)$ and $T_{\ell}$ as in Lemma 7.1. But, since $B_{\ell}=C_{\ell}=0$, we have that $B$, hence $\sigma^{r}\theta_{\ell}$ depends only on $y_{1}$ and $\theta_{\ell}$. Consider now the map
@@
$\theta_{\ell}(Z,W)=e\left(\frac{1}{2}\left(xX_{1}+iyY_{1}\right)\right),$
@@
is independent of $a,b$ and so belongs in the image of $\tau$. But by [deitmar_krieg, Proposition 1.1], $\tau^{-1}\circ R\circ\tau$ is a simple tensor in $\mathbb{D}(\mathbb{H}_{1}\times\mathbb{H}_{1})$, so the problem is reduced in the one-dimensional case and $S_{1}[\ell]=R_{W}[\ell]=0$, i.e., suffices to show
@@
where now we have $\displaystyle{\delta_{k}=\frac{k}{2iy}+\frac{\partial}{\partial z}}$ in the one-dimensional case. But we have
$\delta_{k}^{(r)}y^{\frac{n+2}{2}}=\textup{(const)}\times y^{n/4+1}.$
Also, from [bocherer, page 807], we have that in the one-dimensional case
$R(1,2+n/4)=4y^{2}\frac{\partial}{\partial{z}}\frac{\partial}{\partial{\overline{z}}}-\frac{n}{4}\left(\frac{n}{4}+1\right)$
[…]
Finally, combining the two above results, we obtain the following Lemma.
Lemma 7.3.
Let $k=-n/2$ and $r=n/4$, as before. Given $l\in\mathbb{R}\textup{, }\epsilon>0$, and a compact subset $\mathcal{C}$ of $\mathcal{H}_{S}$, there exists a constant $C>0$ such that
holds for all $W\in\mathcal{C}$ and $Z=X+iY\in\mathbb{H}_{2}$, with $Y\geq\epsilon 1_{2}$.
Proof.
From Proposition 7.2, the fact that $R$ is well-behaved (see (18)), (16) and the relations we have obtained in the proof of Lemma 7.1, we can write
(19) $R\left[\delta_{k}^{(r)}\Theta\right](Z,W)=(\det Y)^{\frac{n+2}{2}}\sum_{\begin{subarray}{c}\ell\in\textup{Mat}_{n+4,2}(\mathbb{Z})\\ \textup{rank }\ell=2\end{subarray}}g(S_{1}[\ell],R_{W}[\ell],Y)\theta_{\ell}(Z,W),$
for some polynomial $g$ in the entries of $S_{1}[\ell],R_{W}[\ell]$, and $Y$. The rest of the proof now follows in the same way as in [deitmar_krieg, Proposition 2.1, (b)]. ∎
@@ 8. Theta Correspondence
In this Section, we finally give the theta correspondence between the Klingen-type Eisenstein series of $\textup{SO}(2,n+2)$ and the Siegel-type Eisenstein series for $\textup{Sp}_{2}$. We start with the following definition.
@@
Let $\chi=\chi_{S}$ denote the character of Proposition 6.5. Let also
$P_{2,0}:=\left\{\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\textup{Sp}_{2}(\mathbb{Z})\mid C=0\right\}$
denote the Siegel parabolic subgroup of $\textup{Sp}_{2}$. Notice that $P_{2,0}\cap\Gamma_{0}^{(2)}(q)=P_{2,0}$. For $s\in\mathbb{C}$ with $\textup{Re}(s)>3/2$, we then define the Siegel Eisenstein series with respect to $P_{2,0}$ and with character $\chi$ as
(20) $\widetilde{E}(Z,\chi,s):=\sum_{\gamma\in P_{2,0}\backslash\Gamma_{0}^{(2)}(q)}\overline{\chi(\gamma)}(\det(\textup{Im}(\gamma Z)))^{s}.$
If $\delta=\begin{pmatrix}A&B\\ 0&D\end{pmatrix}\in P_{2,0}$, we have $\det D=\pm 1$, so $\chi(\delta)=\psi(\det D)=1$ by Proposition 6.5 (recall that $4\mid n$). Therefore, the Eisenstein series is well-defined. This Eisenstein series has a meromorphic continuation to $\mathbb{C}$ (see [kalinin, Theorem 1, (3)]). Finally, using the fact that $\chi$ is a character, we have
@@
for all $\gamma\in\Gamma_{0}^{(2)}(q)$. The main Theorem of the paper can now be stated as follows:
@@
Let $S$ have rank $n$, with $4\mid n$ and be such that $\#\mathcal{C}^{1}(\Gamma_{S})=1$. Let also $k=-n/2$ and $r=n/4$, as before. Define
$\phi_{2}(s):=s\left(s-\frac{1}{2}\right),\textup{ }\xi(s):=\pi^{-s/2}\Gamma(s/2)\zeta(s),$
for $s\in\mathbb{C}$. We then have for $\hbox{Re}(s)>n+1$
@@
where $\displaystyle{\gamma_{S}(s):=\frac{1}{[\textup{Sp}_{2}(\mathbb{Z}):\Gamma_{0}^{(2)}(q)]}(-4)^{-r}\phi_{2}(s/2-2r)\phi_{2}(s/2)\prod_{j=1}^{r}\phi_{2}\left(\frac{s-(2j-1)}{2}\right)}$.
@@
Let
$I:=[\textup{Sp}_{2}(\mathbb{Z}):\Gamma_{0}^{(2)}(q)]\left\langle\widetilde{E}(Z,\chi,(s+1)/2-r),R[\delta_{k}^{(r)}\Theta](Z,W)\right\rangle_{\Gamma_{0}^{(2)}(q)}.$
First of all, this integral is well-defined because of Lemma 7.3. We then have
$\displaystyle I$ $\displaystyle=\mathop{\text{\LARGE$\displaystyle\int_{\text{\normalsize$\displaystyle\scriptstyle\Gamma_{0}^{(2)}(q)\backslash\mathbb{H}_{2}$}}$}}\nolimits\left[\sum_{\gamma\in P_{2,0}\backslash\Gamma_{0}^{(2)}(q)}\overline{\chi(\gamma)}(\det(\textup{Im}(\gamma Z)))^{(s+1)/2-r}\overline{R[\delta_{k}^{(r)}\Theta](Z,W)}\right]\hbox{d}^{*}Z$
$\displaystyle=\mathop{\text{\LARGE$\displaystyle\int_{\text{\normalsize$\displaystyle\scriptstyle\Gamma_{0}^{(2)}(q)\backslash\mathbb{H}_{2}$}}$}}\nolimits\left[\sum_{\gamma\in P_{2,0}\backslash\Gamma_{0}^{(2)}(q)}\overline{\chi(\gamma)}(\det(\textup{Im}(\gamma Z)))^{(s+1)/2-r}\chi(\gamma)\overline{R[\delta_{k}^{(r)}\Theta](\gamma Z,W)}\right]\hbox{d}^{*}Z$